Look Manufacturing Corporation has a traditional costing system in which it applies manufacturing overhead to its products using a predetermined overhead rate based on direct labor-hours (DLHs). The company has two products, N06D and M09K, about which it has provided the following data:

N06D M09K
Direct materials per unit $ 31.50 $ 66.30
Direct labor per unit $ 14.00 $ 26.00
Direct labor-hours per unit 0.20 1.00
Annual production (units) 45,300 21,200

The company's estimated total manufacturing overhead for the year is $1,492,786 and the company's estimated total direct labor-hours for the year is 30,260.

The company is considering using a variation of activity-based costing to determine its unit product costs for external reports. Data for this proposed activity-based costing system appear below:

Activities and Activity Measures Estimated
Overhead Cost
Supporting direct labor (DLHs) $ 665,720
Setting up machines (setups) 445,266
Parts administration (part types) 381,800
Total $ 1,492,786
Expected Activity
N06D M09K Total
DLHs 9,060 21,200 30,260
Setups 1,570 989 2,559
Part types 668 252 920

The manufacturing overhead that would be applied to a unit of product N06D under the company's traditional costing system is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • $4.86

  • $9.87

  • $13.39

  • $21.40
Answer #1

Answer : $9.87

Trading Costing System

Predetermined overhead rate = Estimated Manufacturing OH / Estiamted Direct Labor Hours

Predetermined Overhead Rate = 1,492,786 /30,260 = 49.33

Manufacturing overheads Applied to N06D = 49.33*9060 =446.929.80

Manufactruing overhead Per Unit = 446,929.80 / 45,300 = $9.87 (Answer)